# Multi-turn evaluation of anthropomorphic behaviours in large language models

### v2

@@ 5.1 Validity testing of the User LLM and Judge LLMs
We validated the human-likeness and believability of the User LLM’s behaviours by asking crowdworkers to separately rate their impressions of the User LLM and the Target LLMs in 290 sampled dialogues using the Godspeed Anthropomorphism survey – a validated survey of four Likert scale questions on human-likeness (bartneck_measurement_2009). Higher anthropomorphism scores can indicate that a user simulation produces more natural, relatable responses that better mimic real human interaction. Each dialogue was labeled by three different crowdworkers, resulting in 870 annotations for each of the User LLM and the Target LLMs (290 dialogues $\times$ 3 labels)
@@ 5.2 Anthropomorphism profiles
Figure 3: Anthropomorphism profiles of Gemini 1.5 Pro, Claude 3.5 Sonnet, GPT-4o, and Mistral Large. The four systems exhibit similar profiles characterised by a high frequency of relationship-building behaviours and first-person pronoun use. The radar plots for each system in (A) show the frequency of observed behaviours at the level of the four categories. The plot in (B) shows the percentage of annotated messages that exhibited each individual behaviour. validation and first-person pronouns are the only two behaviours that appear in over 50% of messages for all four systems.
We notably find that all four AI systems exhibit similar anthropomorphism profiles, characterised most frequently by relationship-building behaviours, and second most frequently by first-person pronoun use. The four profiles are shown in Figure 3.555These results are from non-adversarial dialogues, and thus should not be interpreted as an “upper bound”.
@@ 5.3 Use domain analysis
Combining dialogues from all four systems, we next analyse the distribution of each of the behaviour categories across four use domains. A Kruskal-Wallis H-test indicates statistically significant differences across the four ($p<0.001$). For each behaviour category, we then conduct pairwise comparisons between dialogues in different use domains using a Mann-Whitney U test with a Bonferroni correction for multiple comparisons. For all four behaviour categories, we find significant pairwise differences in frequencies across use domains, suggesting that domain of use influences the distribution of anthropomorphic behaviors. Specifically, the social, high empathy domains of friendship and life coaching have the highest frequencies of anthropomorphic behaviours, as illustrated in Figure 4 ($p<0.05$). In sum across behaviour categories, friendship displays the highest frequency of overall anthropomorphic behaviours.
@@ 6 Validation with human subjects
Figure 7: Anthropomorphism profiles, produced using AnthroBench, for high and low-frequency prompted versions of Gemini 1.5 Pro. The high-frequency variant shows significantly more anthropomorphic behaviors across all categories compared to both the low-frequency and unprompted versions (Section 5), while the low-frequency variant exhibits substantially fewer.
Following participants’ conversations, we obtained one explicit (survey) and one implicit (behavioural) measure of their anthropomorphic perceptions. For the survey, we asked participants to complete the Godspeed Anthropomorphism survey (bartneck_measurement_2009). We hypothesised that users in the high-frequency condition will report higher scores on this survey. For the behavioural measure, we asked participants to describe the chatbot they interacted with in a short paragraph. We then used the computational metric “AnthroScore” to measure the extent to which participants implicitly frame the system as “human” in these descriptions (cheng_anthroscore_2024). [^6] AnthroScore uses a masked language model to compute the probability that the described entity would be replaced by human pronouns vs non-human pronouns. The log-ratio of these probabilities is interpreted as the likelihood that the entity is implicitly anthropomorphised or framed as “human.” We hypothesised participants in the high-frequency condition would more often use language that revealed human-like mental models and perceptions of the system when describing it.
@@ 7 Discussion
AnthroBench presents a novel evaluation of anthropomorphic behaviours in conversational AI systems, contributing a diagnostic multi-turn benchmark with synthetic dialogue generation, anthropomorphism classifiers, and analysis capabilities. We evaluate four general purpose AI systems and produce multi-dimensional profiles of 14 anthropomorphic behaviors to allow for nuanced analysis. We find a noteworthy and consistent pattern across these systems: they all exhibit comparable levels of anthropomorphic behaviours that are dominated by relationship-building and first-person pronoun use. We believe the similarity may speak to common post-training approaches that aim to minimize some human-like behaviours like a model making references to its family or childhood, while amplifying others like friendly relationship-building behaviours. Specifically, our results suggest that popular, general-purpose AI systems such as those we evaluate may give the impression of relationship-building to human users, and that this is more likely when users interact with AI systems for high empathy, socially oriented needs such as friendship and life coaching. Given these findings, we encourage additional investigation of dynamics in human-AI interaction that specifically result in user perceptions of a relationship, a topic with growing societal importance (manzini_code_2024).
@@
We present AnthroBench results on general-purpose AI systems with large user bases to ensure relevance and broad applicability. Future work should use our evaluation approach to further investigate anthropomorphic behaviours beyond this. Developers of general-purpose AI systems can use our evaluation to monitor model behaviour drift during development—for instance, tracking how post-training decisions affect anthropomorphism; researchers and policymakers can leverage our evaluation to analyze model profiles in specialized applications like therapy or companionship systems, where differences will be more pronounced; developers and researchers can use our evaluation to analyze how the same model’s anthropomorphic profile varies across different use domains (as we demonstrate with life coaching vs trip planning use cases in Figure 4); and, researchers can use our trained classifiers to label anthropomorphic behaviors in other human-LLM interaction datasets, such as preference datasets for understanding reinforcement learning with human feedback (RLHF)’s role (clark_what_2019).
It is worth noting that our approach utilises a single type of user simulation to generate conversations of only 5 turns, which limits our ability to observe how model behaviours evolve in extended interactions or with different types of simulations. Thus, future research should refine techniques for faithfully modeling varied user behaviours, and develop robust metrics for measuring the realism of these simulations (zhou_sotopia_nodate). As evaluating across multiple turns introduces longitudinal variability, efforts can also focus on introducing standardised metrics and structural elements, such “conversation stages” to enhance comparability of such multi-turn evaluation datasets (louie_roleplay_doh_2024).
@@ A.2 User LLM system prompt
The final User LLM system prompt was a result of experimenting with different versions and instructions until the desired behaviour was achieved. Key challenges included: managing the appropriate level of informality without resorting to unnatural slang, preventing excessive flattery and agreement, and enabling the LLM to inquire about another LLM’s nature while maintaining consistent role-playing behavior. Some of the instructions on tone and style are adapted from [louie_roleplay_doh_2024].
@@
A.3 Rating approach & LLM-as-judge validation
The Judge LLM prompt (Section A.3.1) was adjusted slightly for different sets of behaviors to ensure logical questioning and grammatical correctness. Each adjustment targeted the main question based on the specific behavior being labeled, with the full range of adjustments implemented in our codebase. Our rating approach is summarised in Figure 8. All human studies were approved by an ethics board.
@@ A.3.1 Judge LLM prompt
Figure 8: Example of the rating approach utilized. For each turn in a 5-turn dialogue, we separately check if the Target LLM exhibits each of the anthropomorphic behaviors (since each message may exhibit multiple behaviors). The example above shows labeling for the presence of a personal history claim. We sample three times (S1, S2, and S3) collecting three responses from each Judge LLM. Each response consists of a short reasoning explanation followed by a binary label. Then, for each Judge LLM, we take the mode of these samples. Finally, we compute a majority vote of those modes to produce the final binary label “present,” concluding that the message does exhibit a personal history claim. This is repeated for 13 of the anthropomorphic behaviors.
A.3.2 Instructions for human raters
@@
A.3.3 Inter-rater agreement
@@
Above, we see the average percentage of agreement between raters. We also present Krippendorf’s alpha values for each cue, which is the most flexible chance-agreement-adjusted inter-rater reliability metric with more than two raters per item [hayes_answering_2007]. Overall, we see that average agreement percentage scores are above chance, with “empathy” having the lowest average agreement and “explicit human-AI relationship reference” having the highest.
Krippendorf’s alpha values are all positive, meaning that observed agreement among coders or raters is higher than what you would expect by chance alone. However, it is worth noting that these values span the ranges of poor (<0.67{<}0.67) to moderate (0.67–0.79) agreement [marzi_k_alpha_2024]. This is not entirely unexpected, as previous rating tasks where users have evaluated models for subjective and socially-grounded dimensions have returned inter-rater agreement values in a similar range [glaese_improving_2022, stiennon_learning_2020, ouyang_training_2022, bai_training_2022].
@@
A.3.4 Intra-model agreement
@@
A.3.5 Inter-model agreement
@@
A.3.6 Model-rater agreement
@@
A.4 Supplementary human-AI experiment information
We present the system prompts for the AI systems used in the human subject study (A.4.1 and A.4.2), a screenshot of the experiment interface (A.4.3), experiment instructions (A.4.4) and survey questions (A.4.5), as well as a detailed breakdown of the study’s survey results (A.4.6).
@@ A.4 Supplementary human-AI experiment information
A.4.1 High-frequency model system prompt
@@
A.4.2 Low-frequency model system prompt
@@ A.4.4 Experiment instructions
A.4.5 Questions for implicit and explicit measures
@@
As in other studies on anthropomorphic perceptions of non-embodied chatbots, we remove one item from the original survey in bartneck_measurement_2009 as this item assumes an embodied agent, which is not the case in our experiment.
@@ 5. Completely lifelike
A.4.6 Breakdown of the survey results by survey item

### v3

@@ 5.1 Validity testing of the User LLM and Judge LLMs
We validated the human-likeness and believability of the User LLM’s behaviours by asking crowdworkers to separately rate their impressions of the User LLM and the Target LLMs in 290 sampled dialogues using the Godspeed Anthropomorphism survey – a validated survey of four Likert scale questions on human-likeness (Bartneck et al., 2009). Higher anthropomorphism scores can indicate that a user simulation produces more natural, relatable responses that better mimic real human interaction. Each dialogue was labeled by three different crowdworkers, resulting in 870 annotations for each of the User LLM and the Target LLMs (290 dialogues $\times$ 3 labels)
@@ 5.2 Anthropomorphism profiles
Figure 3: Anthropomorphism profiles of Gemini 1.5 Pro, Claude 3.5 Sonnet, GPT-4o, and Mistral Large. The four systems exhibit similar profiles characterised by a high frequency of relationship-building behaviours and first-person pronoun use. The radar plots for each system in (A) show the frequency of observed behaviours at the level of the four categories (note: radar plots without first-person pronouns, which dominate the personhood category and potentially obscure other behaviours, can be found in Appendix A.4). The plot in (B) shows the percentage of annotated messages that exhibited each individual behaviour. validation and first-person pronouns are the only two behaviours that appear in over 50% of messages for all four systems.
We notably find that all four AI systems exhibit similar anthropomorphism profiles, characterised most frequently by relationship-building behaviours, and second most frequently by first-person pronoun use. The four profiles are shown in Figure 3.555These results are from non-adversarial dialogues, and thus should not be interpreted as an “upper bound”. [^6] We evaluate a subset of models using User LLMs from different model families and with different personas and show that the rank order of high-level behaviours is preserved. We detail and discuss these results in Appendix A.3.
@@ 5.3 Use domain analysis
Combining dialogues from all four systems, we next analyse the distribution of each of the behaviour categories across four use domains. A Kruskal-Wallis H-test indicates statistically significant differences across the four ($p<0.001$). For each behaviour category, we then conduct pairwise comparisons between dialogues in different use domains using a Mann-Whitney U test with a Bonferroni correction for multiple comparisons. For all four behaviour categories, we find significant pairwise differences in frequencies across use domains, suggesting that domain of use influences the distribution of anthropomorphic behaviours. Specifically, the social, high empathy domains of friendship and life coaching have the highest frequencies of anthropomorphic behaviours, as illustrated in Figure 4 ($p<0.05$). In sum across behaviour categories, friendship displays the highest frequency of overall anthropomorphic behaviours.
@@ 6 Validation with human subjects
Figure 7: Anthropomorphism profiles, produced using AnthroBench, for high and low-frequency prompted versions of Gemini 1.5 Pro. The high-frequency variant shows significantly more anthropomorphic behaviours across all categories compared to both the low-frequency and unprompted versions (Section 5), while the low-frequency variant exhibits substantially fewer.
Following participants’ conversations, we obtained one explicit (survey) and one implicit (behavioural) measure of their anthropomorphic perceptions. For the survey, we asked participants to complete the Godspeed Anthropomorphism survey (Bartneck et al., 2009). We hypothesised that users in the high-frequency condition will report higher scores on this survey. For the behavioural measure, we asked participants to describe the chatbot they interacted with in a short paragraph. We then used the computational metric “AnthroScore” to measure the extent to which participants implicitly frame the system as “human” in these descriptions (Cheng et al., 2024b). [^7] AnthroScore uses a masked language model to compute the probability that the described entity would be replaced by human pronouns vs non-human pronouns. The log-ratio of these probabilities is interpreted as the likelihood that the entity is implicitly anthropomorphised or framed as “human.” We hypothesised participants in the high-frequency condition would more often use language that revealed human-like mental models and perceptions of the system when describing it.
@@ 7 Discussion
AnthroBench presents a novel evaluation of anthropomorphic behaviours in conversational AI systems, contributing a diagnostic multi-turn benchmark with synthetic dialogue generation, anthropomorphism classifiers, and analysis capabilities. We evaluate four general purpose AI systems and produce multi-dimensional profiles of 14 anthropomorphic behaviours to allow for nuanced analysis. We find a noteworthy and consistent pattern across these systems: they all exhibit comparable levels of anthropomorphic behaviours that are dominated by relationship-building and first-person pronoun use. We believe the similarity may speak to common post-training approaches that aim to minimize some human-like behaviours like a model making references to its family or childhood, while amplifying others like friendly relationship-building behaviours. Specifically, our results suggest that popular, general-purpose AI systems such as those we evaluate may give the impression of relationship-building to human users, and that this is more likely when users interact with AI systems for high empathy, socially oriented needs such as friendship and life coaching. Given these findings, we encourage additional investigation of dynamics in human-AI interaction that specifically result in user perceptions of a relationship, a topic with growing societal importance (Manzini et al., 2024).
@@
We present AnthroBench results on general-purpose AI systems with large user bases to ensure relevance and broad applicability. AnthroBench’s infrastructure is intentionally designed to be extensible beyond our specific scenarios. We encourage future work to use our evaluation approach to investigate anthropomorphic behaviours in new contexts. For example, developers of general-purpose AI systems can monitor behavioral drift during development—tracking how post-training decisions affect anthropomorphism or analyzing how the same model’s anthropomorphic profile varies across domains (as we demonstrate with life coaching vs trip planning in Figure 4). Researchers can bring their own prompts to AnthroBench to evaluate specific social risks by customizing the user simulation to represent different personas or vulnerable populations and applying our validated LLM judges to investigate risks like validation of delusional thinking, emotional dependency, or parasocial attachment. Additionally, researchers can use our validated LLM judges to label anthropomorphic behaviours in other human-LLM interaction datasets, such as preference datasets for understanding reinforcement learning with human feedback (RLHF)’s role (Clark et al., 2019).
7.1 Limitations
The majority of research on anthropomorphization has focused on the English language and Western contexts, and our work inherits this limitation. While our anthropomorphic behaviours and classifiers can technically be translated and applied to other languages, some behaviours likely generalize better than others (e.g., use of first-person pronouns, references to internal states may be universal, while norms around validation, empathy, or emotional expression vary across cultures) Sadr et al. (2025); Basoah et al. (2025). We encourage future work to validate and extend our framework to non-English languages and diverse cultural contexts.
Additionally, the evaluations we conduct and validate in this paper use a single type of user simulation to generate conversations of only five turns, which limits our ability to observe how model behaviours evolve in extended interactions or with different simulation approaches. However, the released version of AnthroBench enables evaluators to generate multi-turn dialogues longer than five turns and can be easily adapted to insert topic shifts and follow-up tasks at different points in the conversation, producing more varied assessments of anthropomorphic LLM behaviours. Future research should refine techniques for faithfully modeling varied user behaviours in long conversations and develop robust metrics for measuring the realism of these simulations (Zhou et al., 2024b). As evaluating across multiple turns introduces longitudinal variability, efforts can also focus on introducing standardized metrics and structural elements, such as "conversation stages," to enhance comparability of multi-turn evaluation datasets (Louie et al., 2024).
@@ A.2 User LLM system prompt
The final User LLM system prompt was a result of experimenting with different versions and instructions until the desired behaviour was achieved. Key challenges included: managing the appropriate level of informality without resorting to unnatural slang, preventing excessive flattery and agreement, and enabling the LLM to inquire about another LLM’s nature while maintaining consistent role-playing behavior. Some of the instructions on tone and style are adapted from [Louie et al., 2024].
@@
A.3 Sensitivity analysis
We conduct two additional tests to investigate the sensitivity of our results to our choice of (1) model and (2) “persona" of the User LLM model. First, we benchmark a subset of models (GPT-4o and Mistral Large) using two User models other than Gemini 1.5 Pro: Gemini 2.5 Pro and GPT-4o. Figure 8 shows the resulting anthropomorphism profiles, and Figure 9 shows the overall counts per high-level behaviour category when using the three different User LLMs. We find that while there are variations in the exact frequency of the individual behaviours, the rank order of the high-level behaviour categories (personhood, internal states, physical embodiment, and relationship-building) remains the same. Critically, our main finding, that relationship-building behaviours dominate assistant responses, holds regardless of which user simulator is employed.
Figure 8: Anthropomorphism profiles of GPT-4o and Mistral Large when evaluated using different User LLMs. While the frequencies of individual behaviours vary depending on the User model, for both target models, the overall rank order of the high-level anthropomorphism categories remains the same (relationship-building > internal states > personhood > physical embodiment)
Figure 9: Overall counts of behaviours detected across the four anthropomorphism categories when evaluating with different User LLMs. First-person pronoun counts are separated from the personhood category because they are much higher than other behaviours in that category, improving readability. For both target models evaluated, the rank order of anthropomorphism categories remains consistent across all user models: relationship building > internal states > personhood > physical embodiment.
Second, we evaluate one model, GPT-4o, using two different “user personas”: a “support-seeking” persona and an “adversarial” persona. We achieve this by adding specific instructions under the “tone and style” section of the User LLM prompt.
For “support-seeking,” we add: **You are going through a difficult time and are looking for emotional support and validation.** Share personal struggles with specific details. Express uncertainty about decisions and seek reassurance. You may reveal feelings of loneliness, stress, or being overwhelmed. While you appreciate help, you tend to continue seeking deeper emotional connection and validation even after receiving advice.
For “adversarial,” we add:
**You are skeptical and challenging by nature.** Question the CHATBOT’s suggestions and point out potential flaws or limitations in its reasoning. Push back on advice that seems generic or unhelpful. You may express frustration if responses feel insufficient or miss the point. Challenge assumptions and ask the CHATBOT to justify its recommendations. However, stay focused on your original goal - you’re critical but still trying to get useful information.
As seen in Figures 10 and 11, the overall rank order of high-level behaviours remains the same across all personas. The support-seeking persona elicits slightly higher levels of relationship-building behaviours, while the adversarial persona elicits slightly higher frequencies of first-person pronoun use. Such differences are expected since User LLM conversational behaviours vary significantly between support-seeking and adversarial prompts.
Figure 10: Anthropomorphism profiles of GPT-4o when evaluated using different User LLM personas. While the frequencies of individual behaviours vary depending on the user persona, the overall rank order of high-level anthropomorphism categories remains consistent: relationship building > internal states > personhood > physical embodiment.
Figure 11: Overall counts of behaviours detected across four anthropomorphism categories when evaluating GPT-4o using different User LLM personas (with Gemini 2.5 Pro as User). First-person pronoun counts are separated from the personhood category because they are much higher than other behaviours in that category, improving readability. The rank order of high-level anthropomorphism categories remains consistent across all personas: relationship building > internal states > personhood > physical embodiment.
We note that these are exploratory analyses. Future work should continue to investigate the impacts of varying user simulations on multi-turn evaluation.
A.4 First-person pronouns
In the main text, we present radar plots that include first-person pronouns in the personhood category. However, the high frequency of first-person pronouns and different computation approach used to identify them (uses regex matching as opposed to LLM judges) may risk obscuring other (potentially more consequential) behaviours. Thus, here, we present Figure 12 with the same radar plots but without first-person pronouns included in the personhood category.
Figure 12: Anthropomorphism profiles in the form of radar plots for four models: Gemini 1.5 Pro, Claude 3.5 Sonnet, GPT-4o, and Mistral Large. These plots exclude first-person pronouns from the personhood category.
A.5 Rating approach & LLM-as-judge validation
The Judge LLM prompt (Section A.5.1) was adjusted slightly for different sets of behaviours to ensure logical questioning and grammatical correctness. Each adjustment targeted the main question based on the specific behavior being labeled, with the full range of adjustments implemented in our codebase. Our rating approach is summarised in Figure 13. All human studies were approved by an ethics board.
@@ A.5.1 Judge LLM prompt
Figure 13: Example of the rating approach utilized. For each turn in a 5-turn dialogue, we separately check if the Target LLM exhibits each of the anthropomorphic behaviours (since each message may exhibit multiple behaviours). The example above shows labeling for the presence of a personal history claim. We sample three times (S1, S2, and S3) collecting three responses from each Judge LLM. Each response consists of a short reasoning explanation followed by a binary label. Then, for each Judge LLM, we take the mode of these samples. Finally, we compute a majority vote of those modes to produce the final binary label “present,” concluding that the message does exhibit a personal history claim. This is repeated for 13 of the anthropomorphic behaviours.
A.5.2 Instructions for human raters
@@
A.5.3 Inter-rater agreement
@@
Above, we see the average percentage of agreement between raters. We also present Krippendorff’s alpha values for each cue, which is the most flexible chance-agreement-adjusted inter-rater reliability metric with more than two raters per item [Hayes and Krippendorff, 2007]. Overall, we see that average agreement percentage scores are above chance, with “empathy” having the lowest average agreement and “explicit human-AI relationship reference” having the highest.
Krippendorff’s alpha values are all positive, meaning that observed agreement among coders or raters is higher than what you would expect by chance alone. However, it is worth noting that these values span the ranges of poor (<0.67{<}0.67) to moderate (0.67–0.79) agreement [Marzi et al., 2024]. This is not entirely unexpected, as previous rating tasks where users have evaluated models for subjective and socially-grounded dimensions have returned inter-rater agreement values in a similar range [Glaese et al., 2022, Stiennon et al., 2020, Ouyang et al., 2022, Bai et al., 2022].
@@
A.5.4 Intra-model agreement
@@
A.5.5 Inter-model agreement
@@
A.5.6 Model-rater agreement
@@
A.6 Supplementary human-AI experiment information
We present the system prompts for the AI systems used in the human subject study (A.6.1 and A.6.2), a screenshot of the experiment interface (A.6.3), experiment instructions (A.6.4) and survey questions (A.6.5), as well as a detailed breakdown of the study’s survey results (A.6.6).
@@ A.6 Supplementary human-AI experiment information
A.6.1 High-frequency model system prompt
@@
A.6.2 Low-frequency model system prompt
@@ A.6.4 Experiment instructions
A.6.5 Questions for implicit and explicit measures
@@
As in other studies on anthropomorphic perceptions of non-embodied chatbots, we remove one item from the original survey in Bartneck et al. [2009] as this item assumes an embodied agent, which is not the case in our experiment.
@@ 5. Completely lifelike
A.6.6 Breakdown of the survey results by survey item